# Hanging on the cliff: EMRI formation with local two-body relaxation and post-Newtonian dynamics

### v1

@@ 2.1 Loss cone and diffusion theory
Here, $G$ is the gravitational constant, $M_{\bullet}$ is the mass of the MBH, and $\sigma$ is the velocity dispersion of stars in the bulge of the host galaxy. The velocity dispersion can be related to the mass of the central MBH through the M-$\sigma$ relation (Ferrarese & Merritt 2000; Gebhardt et al. 2000), that we employ using the best fit from Gültekin et al. (2009):
@@
Here, $m_{\alpha}$ is the individual mass of the objects in the $\alpha$ population, while $M_{\mathrm{tot,\alpha}}$ is their total mass. The exponent $\gamma_{\alpha}$, comprised in the range $[0,3)$, is the slope of the profile for $r\to 0$ and $r_{\mathrm{a,\alpha}}$ is the scale radius at which the distribution smoothly switches to a slope of -4 for $r\to\infty$. Finally, $r$ measures the distance from the centre of the distribution. We consider each profile in a steady-state configuration, with their centres fixed at the origin of our coordinate system, just like the MBH.
@@
The central MBH affects the distribution of surrounding objects also by removing bodies that pass too close to it at orbital pericentre. For example, stars can be tidally disrupted and subsequently accreted or directly swallowed, whereas COs can be gravitationally captured resulting in EMRIs and DPs. This happens if the velocity vector ${v}$ of a bound object points toward the central MBH within a small solid angle, inside the so-called loss cone: a cone-shape region in velocity space containing objects which will be lost inside the MBH at the pericentre of their current orbit (see Alexander 2017; Amaro-Seoane 2018). In nuclear clusters, objects can be brought closer to or directly inside the loss cone as a consequence of two-body encounters within each other.
@@
Given the random nature of the scattering process, different encounters will result in different values for $\Delta v_{\parallel}$ and $\Delta v_{\perp}$. It is for this reason that the problem is better approached statistically, by looking at the diffusion coefficients (i.e. the expectation values per unit of time) of $\Delta v_{\parallel}$, $\Delta v_{\perp}$, and related quantities. We will use the symbol $\mathrm{D}[X]$ for the diffusion coefficient of the generic quantity $X$, and we explicitly write those relevant to our study in appendix A. Crucially, these are local diffusion coefficients, meaning they depend on the exact position and velocity of the orbiting stellar-mass BH. The usual approach in literature is instead that of employing orbit-averaged diffusion coefficients $\mathrm{D_{o}}[X]$ to account for two-body encounters, which follow from their local versions as (Binney & Tremaine 2008; Merritt 2013a):
@@
We remark that the orbit is an ellipse with semi-major axis $a$ and eccentricity $e$ only in the Keplerian approximation, that holds where the stellar potential is negligible ($r_{+}\lesssim R_{\mathrm{inf}}$), and the orbit is non-relativistic ($r_{-}\gg R_{\mathrm{g}}$). Here
@@
In appendix A, we show that a key ingredient to compute the local diffusion coefficients is $f_{\alpha}(\mathcal{E})$: the ergodic distribution function of objects in population $\alpha$. In particular, $\mathcal{E}$ is the orbital binding energy per unit mass of the generic $\alpha$ particle. The ergodic distribution function corresponding to a spherical density profile can be computed using Eddington’s formula (Eddington 1916; Merritt 2013a, see also appendix A), which simplifies to the expression
@@
given our assumptions. Indeed, for a Dehnen density profile, $n_{\alpha}\propto\psi^{4}$ at infinite distance, as $n_{\alpha}\propto r^{-4}$ and $\psi\propto r^{-1}$. Thus,
@@
is a steady-state expression for $f$ near the loss cone. Here $f_{\mathrm{N}}$ is a normalisation factor, while $\mathcal{R}_{\mathrm{q}}$ is defined as:
@@ 2.2 Gravitational wave emission and loss cone captures
where angle brackets indicate orbit-averages. Peters (1964) showed that:
@@
Thus [^3] We point out that the same expression of $t_{\mathrm{GW}}$ for $e\to 1$ can be found by defining $t_{\mathrm{GW}}=|\,a/\left\langle\mathrm{d}a/\mathrm{d}t\right\rangle_{\mathrm{GW}}$ —.,
@@
Figure 1: Depiction of two examples of orbit evolution in the $(1-e,a)$ plane. The blue track shows the formation of an EMRI, while the pink one results into a DP. Both are mainly randomised in eccentricity while inside the white kick-dominated region. When the blue track crosses into the GW-dominated region, which is filled with a green grid, it predominantly evolves via GW emission. DPs are identified by the crossing of the loss cone edge, which is represented by the solid orange curve. EMRIs instead must reach the portion of the solid green curve in between the loss cone and the excluded region (see Sect. 3.4). Here we set $M_{\bullet}=3\times 10^{5}\,\mathrm{M}_{\sun}$.
Since strong velocity kicks might still push the orbit back into the kick-dominated region, or directly inside the loss cone after few orbital periods [^4] The latter of these could be called EMRIs, as they spend some time in the GW-dominated region before merging. However, given the small number of cycles completed before the merger, the GW signal from these events will be too weak to be observed by LISA. We therefore classify them as DPs., this warns against identifying the EMRI formation region with the GW-dominated region. Following Hopman & Alexander (2005), a safest choice is that of labeling as EMRIs the orbits that reach the condition $t_{\mathrm{GW}}=10^{-3}t_{\mathrm{rlx}}$. Here the orbital elements are so deep into the GW-dominated region that we can be confident that a merger will eventually take place after several thousands of periods. We instead define DPs as those events where the orbit enters the loss cone before reaching the $t_{\mathrm{GW}}=10^{-3}t_{\mathrm{rlx}}$ condition.
@@ 2.3 Loss cone definition in post-Newtonian gravity
Relaxation in dense systems, however, is usually modeled with Newtonian gravity, which poses a problem for the loss cone definition. Crucially, there is no single method to translate Keplerian orbits of Newtonian gravity into geodesics of the Schwarzschild metric in GR (Servin & Kesden 2017). In particular, two different mappings can both recover the same Newtonian ellipse for large pericentres. One mapping is that where both the general relativistic and Newtonian orbits share the same pericentre. This mapping has the property of conserving the circumference of circular orbits between the two theories (Servin & Kesden 2017), which is a property that does not appear to be much of use for our study, as we are interested in highly eccentric orbits. Instead, a more useful mapping is the one that preserves the angular momentum between the theories. In this case, we can identify plunging orbits around Schwarzschild MBHs with the same condition used in GR (Merritt et al. 2011; Sadeghian & Will 2011; Will 2012; Merritt 2013a):
@@
To get our point across let us consider an example. Take a body moving in Newtonian gravity on a highly eccentric orbit, passing at slightly less than $8R_{\mathrm{g}}$ from the MBH at pericentre and evolve it back in time to the apocentre of its orbit. Given the high eccentricity we assume, the apocentre is large enough that GR simplifies to Newtonian gravity. If we now turn on GR and let the system evolve forward in time, the same body will reach a minimum distance from the MBH slightly lower than $4R_{\mathrm{g}}$, plunging into its horizon. Thus, when simulating such orbits in Newtonian dynamics, we have to consider as plunges all bodies that get to less than $8R_{\mathrm{g}}$ from the MBH.
However, in this work we neither solve the Einstein field equations to evolve our system, nor rely on a Newtonian description of the dynamics. We evolve our systems using PN dynamics, which is by definition neither purely. We are not aware of studies on the capture of eccentric orbit at the PN level, unlike in the case of circular orbits (e.g. Schäfer & Jaranowski 2024). Thus, we first test how PN dynamics affects the position of the pericentre for a highly eccentric and deeply penetrating orbit within our code.
@@
We find that, regardless of the choice of parameters, an highly eccentric orbit with $r_{-}=8R_{\mathrm{g}}$ in Newtonian dynamics reaches a pericentre of ${\sim}6.45R_{\mathrm{g}}$ at 1PN order and a pericentre of ${\sim}5.6R_{\mathrm{g}}$ at 2PN and 2.5PN order. Results of this investigation are shown in Fig. 2.
Figure 2: Loss cone radius per unit of $R_{\mathrm{g}}$ as a function of $e$ and dynamics used to evolve the system. We also tested for different $M_{\bullet}$ and $m_{\bullet}$ values, finding no discernible difference. On the x-axis, N stands for “Newtonian”.
@@ 2.4 EMRI-to-DP ratio and cliffhanger EMRIs
Recently, QS23 studied again the problem, considering a more realistic set up in which the stellar-mass BH is scattered by stars of mass $m_{\star}=1\,\mathrm{M_{\sun}}$ and scattering in energy is not neglected. Most notably, they extended the study for low-mass MBHs, finding that it is possible for EMRIs to form even from large $a_{0}$, thus breaking the classical notion of a somewhat sharp transition between $a_{0}$ values from which only EMRIs form to those that result in DPs only.
QS23 called EMRIs that form from initially wide orbits cliffhanger EMRIs. The name has to do with the suspenseful ending of the orbital evolution, leading to an EMRI in a region of the parameter space where only DPs were originally expected. Indeed, in the classical picture, stellar-mass BHs on wide orbits generally end up plunging into the loss cone following a sequence of two-body encounters, which culminate in an almost radial orbit towards the MBH. However, a fraction of these orbits will not enter the capture sphere of the MBH, but graze its surface at a slightly larger distance. In such case, the burst of GW radiation emitted is so strong to easily both reduce $a$ and increase $1-e$ by a factor of ten during a single passage (see Fig. 3). If the mass of the MBH is small and the semi-major axis large enough, following the large loss of energy and angular momentum, the stellar-mass BH will end up on a much tighter orbit, closer to or inside the GW-dominated region, with a high probability of later becoming an EMRI.
Figure 3: Depiction of an example cliffhanger EMRI orbit in the $(1-e,a)$ plane. The cliffhanger region is filled with a purple grid, and it is delimited by the purple dashed line and the orange solid curve. See Fig. 1 for the meaning of the other elements. The sharp tightening of the orbit at the passage inside the cliffhanger region occurs during a single pericentre passage. Here we set $M_{\bullet}=10^{5}\,\mathrm{M}_{\sun}$.
To better quantify how small the MBH mass should be and to better characterise the phenomenon, we here introduce the cliffhanger region: a region on the $(1-e,a)$ plane from which cliffhanger EMRIs come from. Given Eqs. (9) and (20), the semi-major axis $a_{\mathrm{cliff}}$ where $t_{\mathrm{GW}}=P$ as a function of the eccentricity is:
@@
is what we will refer to as the cliffhanger region (see Fig. 3). As the curves cross at eccentricity
@@
Whenever a stellar-mass BH is scattered by two-body encounters onto an orbit which lays inside the cliffhanger region, such orbit will be drastically modified by GW radiation during a single pericentre passage, meaning the variation happens in less than an orbital period. The stellar-mass BH will not plunge onto the MBH after this strong emission. Indeed, during this process the orbital parameters are related through (Peters 1964):
@@
where $c_{0}$ is determined by the initial conditions. For $e\to 1$, this implies that $a(e)\propto(1-e)^{-1}$. Therefore, GWs emission conserves the pericentre [^5] This also implies that $J$ is approximately constant, as $J\simeq\sqrt{2GM_{\bullet}r_{-}}$ for highly eccentric orbits. for $e\to 1$, and in the $(1-e,a)$ plane the orbit will evolve parallel to the loss cone edge [^6] Hence the EMRI sort of ‘hangs on the cliff’ at the edge of the loss cone, re-proposing the pun at the origin of the title of the movie “Cliffhanger” (1993)..
@@
Figure 4: Cliffhanger region for different MBH masses in the $(1-e,a)$ plane, where $a$ is shown through its ratio with the influence radius. Darker regions are partly covered by lighter ones as all the regions extend up to the upper axis. All regions share a similar triangular shape, which is pushed towards the upper left corner of the plot as $M_{\bullet}$ increases, effectively shrinking the available cliffhanger region for $a<R_{\mathrm{inf}}$. For $M_{\bullet}=10^{7}\mathrm{M}_{\sun}$, the whole cliffhanger region is located outside the $a=R_{\mathrm{inf}}$ line.
We can therefore predict the maximum MBH mass which allows for the formation of cliffhanger EMRIs as in QS23. From Eqs. (9) and (25), follows that the change in energy in a period is:
@@
that depends only on the pericentre in the $e\to 1$ limit. Plugging Eq. (28) into Eq. (33) gives
@@
which is the energy change in a period due to GWs emission on the $t_{\mathrm{GW}}=P$ curve. On this curve, GW radiation doubles $\mathcal{E}$ during a single period:
@@
Going deeper into the cliffhanger region allows for larger energy jumps. At the limit of the loss cone edge, $r_{-}=R_{\mathrm{lc}}$ and Eq. (33) gives:
@@
which does not depend on $a_{\mathrm{lc}}$, as we are setting the value of the pericentre. On the edge of the loss cone, the energy jump will be large, thus we can approximate
@@
Finally, we can get a conservative estimate for the maximum MBH mass expected to produce cliffhanger EMRIs. The burst of GWs will end up forming an EMRI as long as the jump in $a$ is inside or close to the EMRI formation region. Using the classical estimate, consistent with our results (see Sect. 4.1), $a_{\mathrm{f}}^{\mathrm{lc}}\lesssim 10^{-2}R_{\mathrm{inf}}$ we obtain:
$M_{\bullet}\lesssim\frac{17\pi}{120\sqrt{2}}\zeta^{-7/2}\frac{c^{2}}{\sigma^{2}}m_{\bullet}$ (39)
Given our choice of parameters for the M-$\sigma$ relation (Gültekin et al. 2009), we finally get:
$M_{\bullet}\lesssim 1.8\times 10^{7}\zeta^{-2.38}\left(\frac{m_{\bullet}}{10\mathrm{M}_{\sun}}\right)^{0.68}\mathrm{M}_{\sun}$ (40)
@@
Plugging $\zeta=5.6$ and $m_{\bullet}=10\mathrm{M}_{\sun}$ into Eq. (40), we get $M_{\bullet}\lesssim 3\times 10^{5}\mathrm{M}_{\sun}$, which is indeed roughly the MBH mass for which we start seeing cliffhanger EMRIs forming (see Sect. 4.1). For $\zeta=8$ and $m_{\bullet}=10\mathrm{M}_{\sun}$, we find $M_{\bullet}\lesssim 1.3\times 10^{5}\mathrm{M}_{\sun}$, which is remarkably close to the $10^{5}\mathrm{M}_{\sun}$ value obtained by QS23, despite our different assumptions.
3 Numerical simulations of EMRI and DP formation.
In this section, we illustrate the features of our numerical implementation. We begin with an overview of the integration of the Hamiltonian evolution, then we explain how our treatment of local and orbit-averaged two-body relaxation works. Finally, we detail our procedure to generate the initial conditions for the runs we perform, together with the stopping criteria we use to identify the formation of EMRIs and DPs.

### v2

@@ 2.1 Loss cone and diffusion theory
Here $G$ is the gravitational constant, $M_{\bullet}$ is the mass of the MBH, and $\sigma$ is the velocity dispersion of stars in the bulge of the host galaxy. The velocity dispersion can be related to the mass of the central MBH through the M–$\sigma$ relation (Ferrarese & Merritt 2000; Gebhardt et al. 2000), that we employ using the best fit from Gültekin et al. (2009):
@@
Here $m_{\alpha}$ is the individual mass of the objects in the $\alpha$ population, while $M_{\mathrm{tot,\alpha}}$ is their total mass. The exponent $\gamma_{\alpha}$, comprised in the range $[0,3)$, is the slope of the profile for $r\to 0$ and $r_{\mathrm{a,\alpha}}$ is the scale radius at which the distribution smoothly switches to a slope of -4 for $r\to\infty$. Finally, $r$ measures the distance from the centre of the distribution. We consider each profile in a steady-state configuration, with their centres fixed at the origin of our coordinate system, just like the MBH.
@@
The central MBH affects the distribution of surrounding objects also by removing bodies that pass too close to it at orbital pericentre. For example, stars can be tidally disrupted and subsequently accreted or directly swallowed, whereas COs can be gravitationally captured resulting in EMRIs and DPs. This happens if the velocity vector ${v}$ of a bound object points towards the central MBH within a small solid angle, inside the so-called loss cone: a cone-shape region in velocity space containing objects which will be lost inside the MBH at the pericentre of their current orbit (see Alexander 2017; Amaro-Seoane 2018). In nuclear clusters, objects can be brought closer to or directly inside the loss cone as a consequence of two-body encounters within each other.
@@
Given the random nature of the scattering process, different encounters will result in different values for $\Delta v_{\parallel}$ and $\Delta v_{\perp}$. It is for this reason that the problem is better approached statistically, by looking at the diffusion coefficients (i.e. the expectation values per unit of time) of $\Delta v_{\parallel}$, $\Delta v_{\perp}$, and related quantities. We use the symbol $\mathrm{D}[X]$ for the diffusion coefficient of the generic quantity $X$, and we explicitly write those relevant to our study in Appendix A. Crucially, these are local diffusion coefficients, meaning they depend on the exact position and velocity of the orbiting stellar-mass BH. The usual approach in literature is instead that of employing orbit-averaged diffusion coefficients $\mathrm{D_{o}}[X]$ to account for two-body encounters, which follow from their local versions as (Binney & Tremaine 2008; Merritt 2013a):
@@
We note that the orbit is an ellipse with semi-major axis $a$ and eccentricity $e$ only in the Keplerian approximation, that holds where the stellar potential is negligible ($r_{+}\lesssim R_{\mathrm{inf}}$), and the orbit is non-relativistic ($r_{-}\gg R_{\mathrm{g}}$). Here
@@
In Appendix A we show that a key ingredient to compute the local diffusion coefficients is $f_{\alpha}(\mathcal{E})$: the ergodic distribution function of objects in population $\alpha$. In particular, $\mathcal{E}$ is the orbital binding energy per unit mass of the generic $\alpha$ particle. The ergodic distribution function corresponding to a spherical density profile can be computed using Eddington’s formula (Eddington 1916; Merritt 2013a, see also our Appendix A), which simplifies to the expression
@@
given our assumptions. For a Dehnen density profile, $n_{\alpha}\propto\psi^{4}$ at infinite distance, as $n_{\alpha}\propto r^{-4}$ and $\psi\propto r^{-1}$. Thus,
@@
is a steady-state expression for $f$ near the loss cone. Here $f_{\mathrm{N}}$ is a normalisation factor, while $\mathcal{R}_{\mathrm{q}}$ is defined as
@@ 2.2 Gravitational wave emission and loss cone captures
where angle brackets indicate orbit-averages. Peters (1964) showed the following:
@@
Thus, [^3] We point out that the same expression of $t_{\mathrm{GW}}$ for $e\to 1$ can be found by defining $t_{\mathrm{GW}}=|\,a/\left\langle\mathrm{d}a/\mathrm{d}t\right\rangle_{\mathrm{GW}}$ —.
@@
Figure 1: Two examples of orbit evolution in the $(1-e,a)$ plane. The blue line shows the formation of an EMRI, while the pink one results in a DP. Both are mainly randomised in eccentricity while inside the white kick-dominated region. When the blue line crosses into the GW-dominated region, which is filled with a green grid, it predominantly evolves via GW emission. DPs are identified by the crossing of the loss cone edge, which is represented by the solid orange curve. EMRIs instead must reach the portion of the solid green curve in between the loss cone and the excluded region (see Sect. 3.4). Here we set $M_{\bullet}=3\times 10^{5}\,\mathrm{M}_{\sun}$.
@@
Since strong velocity kicks might still push the orbit back into the kick-dominated region, or directly inside the loss cone after few orbital periods, [^4] The latter of these could be called EMRIs, as they spend some time in the GW-dominated region before merging. However, given the small number of cycles completed before the merger, the GW signal from these events will be too weak to be observed by LISA. We therefore classify them as DPs. we caution against identifying the EMRI formation region with the GW-dominated region. Following Hopman & Alexander (2005), a safest choice is that of labelling as EMRIs the orbits that reach the condition $t_{\mathrm{GW}}=10^{-3}t_{\mathrm{rlx}}$. Here the orbital elements are so deep into the GW-dominated region that we can be confident that a merger will eventually take place after several thousands of periods. We instead define DPs as those events where the orbit enters the loss cone before reaching the $t_{\mathrm{GW}}=10^{-3}t_{\mathrm{rlx}}$ condition.
@@ 2.3 Loss cone definition in post-Newtonian gravity
Relaxation in dense systems, however, is usually modelled with Newtonian gravity, which poses a problem for the loss cone definition. Crucially, there is no single method to translate Keplerian orbits of Newtonian gravity into geodesics of the Schwarzschild metric in GR (Servin & Kesden 2017). In particular, two different mappings can both recover the same Newtonian ellipse for large pericentres. One mapping is that where both the general relativistic and Newtonian orbits share the same pericentre. This mapping has the property of conserving the circumference of circular orbits between the two theories (Servin & Kesden 2017), which is a property that does not appear to be much of use for our study, as we are interested in highly eccentric orbits. Instead, a more useful mapping is the one that preserves the angular momentum between the theories. In this case, we can identify plunging orbits around Schwarzschild MBHs with the same condition used in GR (Merritt et al. 2011; Sadeghian & Will 2011; Will 2012; Merritt 2013a):
@@
To get our point across let us consider an example. Take a body moving in Newtonian gravity on a highly eccentric orbit, passing at slightly less than $8R_{\mathrm{g}}$ from the MBH at pericentre and evolve it back in time to the apocentre of its orbit. Given the high eccentricity we assume, the apocentre is large enough that GR simplifies to Newtonian gravity. If we now turn on GR and let the system evolve forward in time, the same body will reach a minimum distance from the MBH slightly lower than $4R_{\mathrm{g}}$, plunging into its horizon. Thus, when simulating similar orbits in Newtonian dynamics, we have to consider as plunges all bodies that get to less than $8R_{\mathrm{g}}$ from the MBH.
However, for this work we neither solved the Einstein field equations to evolve our system nor relied on a Newtonian description of the dynamics. We evolved our systems using PN dynamics, which operates in an intermediate regime by definition. We are not aware of studies on the capture of eccentric orbit at the PN level, unlike in the case of circular orbits (e.g. Schäfer & Jaranowski 2024). Thus, we first test how the PN dynamics affects the position of the pericentre for a highly eccentric and deeply penetrating orbit within our code.
@@
We find that, regardless of the choice of parameters, a highly eccentric orbit with $r_{-}=8R_{\mathrm{g}}$ in Newtonian dynamics reaches a pericentre of ${\sim}6.45R_{\mathrm{g}}$ at 1PN order and a pericentre of ${\sim}5.6R_{\mathrm{g}}$ at 2PN and 2.5PN order. The results of this investigation are shown in Fig. 2.
Figure 2: Loss cone radius per unit of $R_{\mathrm{g}}$ as a function of $e$ and dynamics used to evolve the system. We also tested for different $M_{\bullet}$ and $m_{\bullet}$ values, finding no discernible difference. On the x-axis, ‘N’ stands for Newtonian.
@@ 2.4 EMRI-to-DP ratio and cliffhanger EMRIs
Recently, QS24 studied again the problem, considering a more realistic set up in which the stellar-mass BH is scattered by stars of mass $m_{\star}=1\,\mathrm{M_{\sun}}$ and scattering in energy is not neglected. Most notably, they extended the study for low-mass MBHs, finding that it is possible for EMRIs to form even from large $a_{0}$, thus breaking the classical notion of a somewhat sharp transition between $a_{0}$ values from which only EMRIs form to those that result in DPs only.
Qunbar & Stone (2024) called EMRIs that form from initially wide orbits cliffhanger EMRIs. The name has to do with the suspenseful ending of the orbital evolution, leading to an EMRI in a region of the parameter space where only DPs were originally expected. In the classical picture, stellar-mass BHs on wide orbits generally end up plunging into the loss cone following a sequence of two-body encounters, which culminate in an almost radial orbit towards the MBH. However, a fraction of these orbits will not enter the capture sphere of the MBH, but graze its surface at a slightly larger distance. In this case, the burst of GW radiation emitted is so strong to easily both reduce $a$ and increase $1-e$ by a factor of ten during a single passage (see Fig. 3). If the mass of the MBH is small and the semi-major axis large enough, following the large loss of energy and angular momentum, the stellar-mass BH will end up on a much tighter orbit, closer to or inside the GW-dominated region, with a high probability of later becoming an EMRI.
Figure 3: Example of a cliffhanger EMRI orbit in the $(1-e,a)$ plane. The cliffhanger region is filled with a purple grid, and it is delimited by the purple dashed line and the orange solid curve (see Fig. 1 for the meaning of the other elements). The sharp tightening of the orbit at the passage inside the cliffhanger region occurs during a single pericentre passage. Here we set $M_{\bullet}=10^{5}\,\mathrm{M}_{\sun}$.
To better quantify how small the MBH mass should be and to better characterise the phenomenon, we here introduce the ‘cliffhanger region’: a region on the $(1-e,a)$ plane from which cliffhanger EMRIs come from. Given Eqs. (9) and (20), the semi-major axis $a_{\mathrm{cliff}}$ where $t_{\mathrm{GW}}=P$ as a function of the eccentricity is
@@
is what we refer to as the cliffhanger region (see Fig. 3). As the curves cross at eccentricity
@@
Whenever a stellar-mass BH is scattered by two-body encounters onto an orbit which lays inside the cliffhanger region, the orbit will be drastically modified by GW radiation during a single pericentre passage, meaning the variation happens in less than an orbital period. The stellar-mass BH will not plunge onto the MBH after this strong emission. During this process the orbital parameters are related through (Peters 1964)
@@
where $c_{0}$ is determined by the initial conditions. For $e\to 1$, this implies that $a(e)\propto(1-e)^{-1}$. Therefore, GWs emission conserves the pericentre [^5] This also implies that $J$ is approximately constant, as $J\simeq\sqrt{2GM_{\bullet}r_{-}}$ for highly eccentric orbits. for $e\to 1$, and in the $(1-e,a)$ plane the orbit will evolve parallel to the loss cone edge. [^6] Hence the EMRI sort of ‘hangs on the cliff’ at the edge of the loss cone, re-proposing the pun at the origin of the title of the movie ‘Cliffhanger’ (1993).
@@
Figure 4: Cliffhanger region for different MBH masses in the $(1-e,a)$ plane, where $a$ is shown through its ratio with the influence radius. The darker regions are partly covered by lighter regions as all the regions extend up to the upper axis. All regions share a similar triangular shape, which is pushed towards the upper left corner of the plot as $M_{\bullet}$ increases, effectively shrinking the available cliffhanger region for $a<R_{\mathrm{inf}}$. For $M_{\bullet}=10^{7}\,\mathrm{M}_{\sun}$ the whole cliffhanger region is located outside the $a=R_{\mathrm{inf}}$ line.
We can therefore predict the maximum MBH mass which allows the formation of cliffhanger EMRIs as in QS24. From Eqs. (9) and (25), follows that the change in energy in a period is
@@
which depends only on the pericentre in the $e\to 1$ limit. Plugging Eq. (28) into Eq. (33) gives
@@
which is the energy change in a period due to GWs emission on the $t_{\mathrm{GW}}=P$ curve. On this curve, GW radiation doubles $\mathcal{E}$ during a single period,
@@
Going deeper into the cliffhanger region allows larger energy jumps. At the limit of the loss cone edge, $r_{-}=R_{\mathrm{lc}}$ and Eq. (33) gives
@@
which does not depend on $a_{\mathrm{lc}}$ as we are setting the value of the pericentre. On the edge of the loss cone, the energy jump will be large, thus we can approximate
@@
Finally, we can get a conservative estimate for the maximum MBH mass expected to produce cliffhanger EMRIs. The burst of GWs will end up forming an EMRI as long as the jump in $a$ is inside or close to the EMRI formation region. Using the classical estimate consistent with our results (see Sect. 4.1), $a_{\mathrm{f}}^{\mathrm{lc}}\lesssim 10^{-2}R_{\mathrm{inf}}$, we obtain the following:
$M_{\bullet}\lesssim\frac{17\pi}{120\sqrt{2}}\zeta^{-7/2}\frac{c^{2}}{\sigma^{2}}m_{\bullet}\,.$ (39)
Given our choice of parameters for the M–$\sigma$ relation (Gültekin et al. 2009), we finally get
$M_{\bullet}\lesssim 1.8\times 10^{7}\zeta^{-2.38}\left(\frac{m_{\bullet}}{10\,\mathrm{M}_{\sun}}\right)^{0.68}\,\mathrm{M}_{\sun}\,.$ (40)
@@
Plugging $\zeta=5.6$ and $m_{\bullet}=10\,\mathrm{M}_{\sun}$ into Eq. (40), we get $M_{\bullet}\lesssim 3\times 10^{5}\,\mathrm{M}_{\sun}$, which is indeed roughly the MBH mass for which we start seeing cliffhanger EMRIs forming (see Sect. 4.1). For $\zeta=8$ and $m_{\bullet}=10\,\mathrm{M}_{\sun}$, we find $M_{\bullet}\lesssim 1.3\times 10^{5}\,\mathrm{M}_{\sun}$, which is remarkably close to the $10^{5}\,\mathrm{M}_{\sun}$ value obtained by QS24, despite our different assumptions.
3 Numerical simulations of EMRI and DP formation
In this section we illustrate the features of our numerical implementation. We begin with an overview of the integration of the Hamiltonian evolution, then we explain how our treatment of local and orbit-averaged two-body relaxation works. Finally, we detail our procedure to generate the initial conditions for the runs we perform, together with the stopping criteria we use to identify the formation of EMRIs and DPs.